# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

### v2

@@ Abstract
We introduce EvoXplain, a simple diagnostic framework that measures the stability of model explanations across repeated training. Rather than analysing the explanation of a single trained model, EvoXplain treats explanations as samples drawn from the stochastic optimisation process itself—without aggregating predictions or constructing ensembles—and examines whether these samples form a single coherent explanation or instead separate into multiple, distinct explanatory modes.
We evaluate EvoXplain on the Breast Cancer and COMPAS datasets across two widely used model classes: Logistic Regression and tree-based models (Random Forests). Although all models achieve high predictive accuracy, their explanations frequently exhibit clear multimodality. Even models commonly assumed to be stable—such as Logistic Regression—can produce multiple, well-separated explanatory basins under repeated training on the same data split. Importantly, these differences are not explained by hyperparameter variation or simple performance trade-offs.
@@ 1 Introduction
Machine learning models are increasingly deployed in domains where explanations matter, including healthcare, public policy, and scientific discovery. Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining. This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (Lipton (2018); Rudin (2019)).
@@
There have been sustained challenges to this assumption in the literature. Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (Breiman (2001); Marx et al. (2020)). Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving optimisation procedures to select arbitrarily among them (D’Amour et al. (2022)). Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (Kindermans et al. (2017); Nogueira et al. (2018); Kumar et al. (2020); Frye et al. (2020)). Collectively, these results suggest that explanation stability cannot be assumed a priori. However, existing work does not provide a practical way to quantify how much explanatory variability exists in a given setting, nor whether such variability is structured or merely diffuse.
@@
To address this gap, we propose EvoXplain, a simple and easy-to-apply framework for assessing the stability of model explanations across repeated training. Rather than evaluating a single attribution vector, EvoXplain treats explanations as outcomes of stochastic optimisation and analyses them as a distribution induced by the training process itself. By clustering explanation samples obtained from repeated instantiations of the same model class—without aggregating predictions or constructing ensembles—EvoXplain determines whether training consistently yields a single explanatory mode or instead produces multiple, distinct explanatory regimes.
@@
EvoXplain does not aim to decide which explanation is correct. Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular optimisation trajectory taken during training. This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (Arrieta and others (2020); Selbst and Barocas (2018); European Commission (2023)). By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
@@ 2.1 Model Performance and Interpretability
A large proportion of studies in machine learning assess models primarily through the lens of accuracy. In the real world, interpretability is often categorized based upon the model being used. Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved. However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (Lipton (2018)). EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
@@ 2.2 Rashomon Effects and Predictive Multiplicity
Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (Breiman (2001); Semengue et al. (2019)). Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (Umbach et al. (2023); Marx et al. (2020); Fisher et al. (2019)), showing that the same risk scores can be justified through different models (Umbach et al. (2023)). Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
@@ 2.4 Feature Importance Instability
Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (Kindermans et al. (2017); Nogueira et al. (2018); Kumar et al. (2020); Agarwal et al. (2021)). Most of these studies focus on instability in some specific values. EvoXplain extends local noise analysis by detecting whether instability forms structured groups of explanations and how often each group appears.
@@ 2.5 Rashomon-Set Explanations in AutoML
A related line of work incorporates model multiplicity into post-hoc explanations by operating over Rashomon sets—collections of near-optimal models within a tolerance of the best model’s loss. Cavus et al. (2025) propose the Rashomon Partial Dependence Profile (Rashomon PDP), which aggregates partial dependence profiles across the Rashomon set and reports uncertainty bands to highlight regions of disagreement. While such approaches usefully acknowledge that multiple near-optimal models exist, they typically treat multiplicity as uncertainty around a single aggregated explanation. EvoXplain studies a different phenomenon: structured, discrete explanatory modes (basins) among equally accurate models, and shows that explanation aggregation can yield a ghost explanation that may not correspond to any realised model instance. In other words, rather than summarising variability around one explanation, EvoXplain detects and quantifies distinct mechanistic regimes within the same training pipeline.
@@ 2.6 Mechanistic Interpretability in Deep Models
The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (Olah et al. (2020)). Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs. On the other hand, EvoXplain focuses on using generic attribution outputs to determine mechanism-level differences.
@@ 2.7 Explainable AI for Science and Metafeature-Space Diagnostics
Recent work in explainable AI for scientific workflows studies where a fixed model tends to improve or degrade, by searching for subgroups of the feature space associated with changes in predictive behaviour. Soares et al. introduce Meta Subspace Analysis (MSA), which discovers exceptional subspaces and relates their metafeatures (e.g., distributional and dependency summaries) to model (mis)behavior via association rules (Soares et al. (2025)). This perspective is complementary to EvoXplain. MSA characterises heterogeneity across data subspaces for a given trained model, whereas EvoXplain characterises heterogeneity across trained model instances under the same pipeline. Thus, MSA does not overlap with EvoXplain’s core object of study (explanation basins over retraining), but motivates future combinations, such as measuring explanation multiplicity within subspaces identified as behaviourally exceptional.
@@ 2.8 Explanation Stability in Safety-Critical Machine Learning
Beyond interpretability research, the safety assurance community has emphasised the need for explicit evidence about model robustness under retraining. Habli and colleagues propose structured assurance methodologies for machine learning components, such as AMLAS and the “BIG Argument” for AI safety (Habli et al. (2025); Hawkins et al. (2021)), which require that system behaviour be justified and verified beyond predictive accuracy alone. EvoXplain contributes to this assurance perspective by offering a concrete diagnostic for explanation stability. When explanations diverge across retrainings of equally accurate models, this reflects a form of epistemic uncertainty that must be accounted for in any rigorous ML safety case. In this sense, EvoXplain does not replace assurance frameworks, but provides a measurable source of evidence that can be incorporated within them.
@@ 3 Problem Formulation
We use interpretability methods to understand how models make decisions, but most analyses implicitly assume that a model has a single, stable explanation once trained. This assumption is not always justified. Due to randomness in the training process, model inductive biases, and model flexibility, it is possible for multiple different explanations to achieve similar predictive performance. A first step toward trustworthy interpretability is therefore to verify whether explanatory multiplicity exists at all.
Formally, consider a model class trained on a fixed dataset . Standard interpretability workflows examine a single trained instance and its corresponding attribution vector . In contrast, this work asks the following question: when is trained multiple times with independent random initializations, do the resulting attribution vectors converge to a single explanatory pattern, or do they separate into multiple distinct patterns?
Let denote independent training runs on , and let denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of to determine:
@@
In practice, we repeat this analysis across multiple train–test splits of to assess the robustness of explanatory behaviour under data resampling. However, the core object of interest remains the distribution of explanations induced by repeated training within a fixed data realisation.
EvoXplain offers an empirical method for answering these questions. Rather than assuming explanations are stable, it treats them as random samples generated by the training process and assesses their structure using clustering and entropy. This reframes interpretability as an empirical question: whether a model class tends to produce a single coherent explanation or multiple competing explanatory mechanisms under repeated instantiation.
@@
Second, we introduce EvoXplain, a simple diagnostic framework that determines if explanations cluster into a single pattern or if they are dispersed across multiple clusters of different patterns. It also measures the occurrence frequency of each distinct explanatory pattern.
@@ 4.1 Repeated Training and Attribution Sampling
Let denote a model class (e.g., Logistic Regression or Random Forests). For each run , we train from an independent random seed while keeping the data split, preprocessing, and hyperparameters fixed. After each training run, we compute feature-level attributions using a chosen interpretability method (here, SHAP), yielding an attribution vector over the input features.
@@
and treats them as samples drawn from the explanation distribution induced by stochastic training. Importantly, EvoXplain does not aggregate predictions or construct ensembles; it analyses explanation variability arising solely from repeated instantiation of the same model class.
@@ 4.3 Clustering and Basin Structure
To assess whether explanations converge to a single pattern or separate into multiple patterns, we apply -means clustering (chosen for simplicity and interpretability) for values of from to . For each , we compute the silhouette score and select
If , explanations concentrate around a single explanatory mode. If , EvoXplain identifies multiple explanatory basins corresponding to distinct explanation patterns reached across training runs.
@@ 4.7 Algorithm Summary
Clustering in explanation space is used here as a diagnostic tool rather than as a claim about the true number of underlying mechanisms. We adopt k-means with silhouette-based selection for simplicity and transparency, and because the resulting clusters are visually apparent in low-dimensional projections and consistent across independent train–test splits. Importantly, the presence or absence of explanation multiplicity does not hinge on fine-grained cluster boundaries: when training converges to a single explanatory regime (e.g., under fixed regularisation), explanations collapse to a single dense region and silhouette analysis consistently selects k = 1, whereas when multiple explanatory regimes are admissible, well-separated modes emerge with stable support across runs and splits. EvoXplain therefore does not rely on clustering to create structure, but uses clustering to quantify structure that is already present in the explanation distribution.
@@ 4.8 Theoretical Interpretation
EvoXplain can be understood through optimisation non-identifiability. Let denote the parameter space of a model class and its loss surface. In high-dimensional or weakly constrained settings, training may converge to multiple parameter configurations with near-optimal loss,
which we refer to as a mechanism manifold. Different training runs may select different points in and yield distinct attribution patterns even when predictive performance is similar.
@@
EvoXplain does not claim these basins correspond to true causal mechanisms. Rather, it demonstrates that optimisation and model capacity alone can induce multiple, stable explanatory behaviours.
@@ 5 Scope and Limitations
EvoXplain does not assert that the clusters it identifies correspond to true causal mechanisms, nor that particular centroids represent any form of ground truth. Its outputs are contingent upon the attribution method used, and the observed structure may also reflect noise or inductive biases inherent to the model or the explanation technique. EvoXplain does not attempt to resolve these questions. Rather, its purpose is to bring such issues to the forefront and to encourage more cautious interpretation of explanation outputs.
EvoXplain has been designed to assess how models behave across repeated training runs. It determines whether a model settles into a single cluster of explanations or whether different explanations emerge across retraining. This perspective comes with several limitations.

### v3

@@ Abstract
We introduce EvoXplain, a simple diagnostic framework that measures the stability of model explanations across repeated training. Rather than analysing the explanation of a single trained model, EvoXplain treats explanations as samples drawn from the training and model selection pipeline itself—without aggregating predictions or constructing ensembles—and examines whether these samples form a single coherent explanation or instead separate into structured regions of explanation space corresponding to distinct explanatory modes.
We evaluate EvoXplain on the Breast Cancer and COMPAS datasets across two widely used model classes: Logistic Regression and tree-based models (Random Forests). Although all models achieve high predictive accuracy, their explanations frequently exhibit clear multimodality. Even models commonly assumed to be stable—such as Logistic Regression—can give rise to structured, low-dimensional explanation manifolds that separate into recurrent regions under repeated training on the same data split. Crucially, these distinct explanation modes coexist even at near-identical hyperparameter configurations, indicating genuine explanation non-identifiability rather than smooth sensitivity to regularisation strength.
@@ 1 Introduction
Machine learning models are increasingly deployed in domains where explanations matter, including healthcare, public policy, and scientific discovery. Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining. This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (Lipton, 2018; Rudin, 2019).
@@
There have been sustained challenges to this assumption in the literature. Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (Breiman, 2001; Marx et al., 2020). Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving training and model selection procedures to select arbitrarily among them (D’Amour et al., 2022). Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (Kindermans et al., 2017; Nogueira et al., 2018; Kumar et al., 2020; Frye et al., 2020). Collectively, these results suggest that explanation stability cannot be assumed a priori. However, existing work does not provide a practical way to quantify how much explanatory variability exists in a given setting, nor whether such variability is structured or merely diffuse.
@@
To address this gap, we propose EvoXplain, a simple and easy-to-apply framework for assessing the stability of model explanations across repeated training. Rather than evaluating a single attribution vector, EvoXplain treats explanations as outcomes of repeated training and model selection and analyses them as a distribution induced by the training pipeline itself. By clustering explanation samples obtained from repeated instantiations of the same model class—without aggregating predictions or constructing ensembles—EvoXplain determines whether training consistently yields a single explanatory mode or instead produces multiple, distinct explanatory regimes.
@@
EvoXplain does not aim to decide which explanation is correct. Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular training and model selection trajectory taken during deployment and retraining. This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (Arrieta and others, 2020; Selbst and Barocas, 2018; European Commission, 2023). By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
@@ 2.1 Model Performance and Interpretability
A large proportion of studies in machine learning assess models primarily through the lens of accuracy. In the real world, interpretability is often categorized based upon the model being used. Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved. However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (Lipton, 2018). EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
@@ 2.2 Rashomon Effects and Predictive Multiplicity
Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (Breiman, 2001; Semengue et al., 2019). Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (Umbach et al., 2023; Marx et al., 2020; Fisher et al., 2019), showing that the same risk scores can be justified through different models (Umbach et al., 2023). Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
@@ 2.4 Feature Importance Instability
Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (Kindermans et al., 2017; Nogueira et al., 2018; Kumar et al., 2020; Agarwal et al., 2021). Most of these studies focus on instability in some specific values. EvoXplain extends local noise analysis by detecting whether instability forms structured groups of explanations and how often each group appears.
@@ 2.5 Rashomon-Set Explanations in AutoML
A related line of work incorporates model multiplicity into post-hoc explanations by operating over Rashomon sets—collections of near-optimal models within a tolerance of the best model’s loss. Cavus et al. (2025) propose the Rashomon Partial Dependence Profile (Rashomon PDP), which aggregates partial dependence profiles across the Rashomon set and reports uncertainty bands to highlight regions of disagreement. While such approaches usefully acknowledge that multiple near-optimal models exist, they typically treat multiplicity as uncertainty around a single aggregated explanation. EvoXplain studies a different phenomenon: low-dimensional explanation manifolds with structured regions among equally accurate models, and shows that explanation aggregation can yield a ghost explanation that may not correspond to any realised model instance. In other words, rather than summarising variability around one explanation, EvoXplain detects and quantifies distinct mechanistic regimes within the same training pipeline.
@@ 2.6 Mechanistic Interpretability in Deep Models
The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (Olah et al., 2020). Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs. On the other hand, EvoXplain focuses on using generic attribution outputs to determine mechanism-level differences.
@@ 2.7 Explainable AI for Science and Metafeature-Space Diagnostics
Recent work in explainable AI for scientific workflows studies where a fixed model tends to improve or degrade, by searching for subgroups of the feature space associated with changes in predictive behaviour. Soares et al. introduce Meta Subspace Analysis (MSA), which discovers exceptional subspaces and relates their metafeatures (e.g., distributional and dependency summaries) to model (mis)behavior via association rules (Soares et al., 2025). This perspective is complementary to EvoXplain. MSA characterises heterogeneity across data subspaces for a given trained model, whereas EvoXplain characterises heterogeneity across trained model instances under the same pipeline. Thus, MSA does not overlap with EvoXplain’s core object of study (explanation basins over retraining), but motivates future combinations, such as measuring explanation multiplicity within subspaces identified as behaviourally exceptional.
@@ 2.8 Explanation Stability in Safety-Critical Machine Learning
Beyond interpretability research, the safety assurance community has emphasised the need for explicit evidence about model robustness under retraining. Habli and colleagues propose structured assurance methodologies for machine learning components, such as AMLAS and the “BIG Argument” for AI safety (Habli et al., 2025; Hawkins et al., 2021), which require that system behaviour be justified and verified beyond predictive accuracy alone. EvoXplain contributes to this assurance perspective by offering a concrete diagnostic for explanation stability. When explanations diverge across retrainings of equally accurate models, this reflects a form of epistemic uncertainty that must be accounted for in any rigorous ML safety case. In this sense, EvoXplain does not replace assurance frameworks, but provides a measurable source of evidence that can be incorporated within them.
@@ 3 Problem Formulation
We use interpretability methods to understand how models make decisions, but most analyses implicitly assume that a model has a single, stable explanation once trained. This assumption is not always justified. Due to randomness in the training process, model inductive biases, and the non-identifiability of admissible training configurations within a modelling pipeline, it is possible for multiple different explanations to achieve similar predictive performance. A first step toward trustworthy interpretability is therefore to verify whether explanatory multiplicity exists at all.
Formally, consider a model class trained on a fixed dataset using a fixed training pipeline (including preprocessing, model family, and admissible hyperparameter configurations). Standard interpretability workflows examine a single trained instance and its corresponding attribution vector . In contrast, this work asks the following question: when is instantiated multiple times under the same pipeline—via different random seeds, bootstrap samples, or admissible hyperparameter choices—do the resulting attribution vectors converge to a single explanatory pattern, or do they separate into multiple distinct patterns?
Let denote independent model instances produced by this pipeline on , and let denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of to determine:
@@
In practice, we repeat this analysis across multiple train–test splits of to assess the robustness of explanatory behaviour under data resampling. However, the core object of interest remains the distribution of explanations induced by repeated admissible instantiations of the same training pipeline on a fixed data realisation.
EvoXplain offers an empirical method for answering these questions. Rather than assuming explanations are stable, it treats them as samples generated by the training pipeline and assesses their structure using clustering and entropy. This reframes interpretability as an empirical question: whether a model class, under realistic retraining and model selection, tends to produce a single coherent explanation or multiple competing explanatory mechanisms.
@@
Second, we introduce EvoXplain, a simple diagnostic framework that determines if explanations cluster into a single pattern or if they are dispersed across multiple basins of different patterns. It also measures the occurrence frequency of each distinct explanatory pattern.
@@ 4.1 Repeated Training and Attribution Sampling
Let denote a model class (e.g., Logistic Regression or Random Forests). For each run , we train under an independently instantiated training pipeline, while keeping the data split and preprocessing fixed. A pipeline instantiation may involve stochastic sources (e.g., data resampling or optimiser initialisation) and/or a sampled hyperparameter configuration, depending on the model class.
After each training run, we compute feature-level attributions using a chosen interpretability method (here, SHAP), yielding an attribution vector over the input features.
@@
and treats them as samples drawn from the explanation distribution induced by repeated instantiation of the training pipeline. Importantly, EvoXplain does not aggregate predictions or construct ensembles; it analyses explanation variability arising solely from admissible variation within a single model class and data split.
For deterministic solvers such as Logistic Regression with lbfgs, run-to-run variation arises from differences in admissible hyperparameter configurations rather than optimiser stochasticity.
4.2 Orthogonal Synthetic Control Dataset
To assess whether explanation multiplicity can arise purely from feature correlation or data redundancy, we introduce a synthetic control dataset with strictly orthogonal input features and a known linear signal structure. Features are generated to be mutually uncorrelated, and labels are assigned using a fixed ground-truth linear rule with additive noise. This setting represents a best-case scenario for explanation stability, where neither feature correlation nor redundant predictors can induce alternative attribution patterns.
We apply the same repeated training and explanation extraction protocol used for real datasets. Under fixed regularisation strength, explanations collapse to a single explanatory regime (). When admissible training configurations are expanded to include a range of regularisation strengths, structured explanation multiplicity re-emerges despite strict feature orthogonality. This confirms that the observed multiplicity is not driven by feature correlation or attribution artefacts, but arises from non-identifiability in the admissible training pipeline itself.
@@ 4.4 Clustering and Basin Structure
To assess whether explanations converge to a single pattern or separate into multiple patterns, we evaluate candidate multi-basin solutions using -means clustering for values of . For each , we compute the silhouette score . We select
subject to the constraint , where is a predefined silhouette threshold. If no candidate satisfies this criterion, the explanation distribution is treated as unimodal and we set .
If this criterion is not met, explanations are treated as forming a single explanatory mode, even if under unconstrained maximisation. When the criterion is satisfied, EvoXplain identifies multiple regions of concentration in explanation space, which we refer to as explanatory basins for diagnostic convenience.
We verified that conclusions are qualitatively stable for reasonable variations of (See Appendix A).
@@ 4.8 Algorithm Summary
Clustering in explanation space is used here as a diagnostic tool rather than as a claim about the true number of underlying mechanisms. We adopt k-means with silhouette-based selection for simplicity and transparency, and because the resulting basin structure is visually apparent in low-dimensional projections and reproducible across independent train–test splits, even when the number or balance of basins varies. Importantly, the presence or absence of explanation multiplicity does not hinge on fine-grained basin boundaries: when training converges to a single explanatory regime (e.g., under fixed regularisation), explanations collapse to a single dense region and silhouette analysis consistently selects k = 1, whereas when multiple explanatory regimes are admissible, well-separated modes emerge with stable support across runs and splits. EvoXplain therefore does not rely on clustering to create structure, but uses clustering to quantify structure that is already present in the explanation distribution.
@@ 4.9 Theoretical Interpretation
EvoXplain can be understood through training-pipeline non-identifiability, which includes but is not limited to optimisation non-identifiability. Let denote the parameter space of a model class and its loss surface. In high-dimensional or weakly constrained settings, or when the training pipeline admits multiple admissible configurations, training may converge to multiple parameter configurations with near-optimal loss,
which we refer to as a mechanism manifold. Different training instantiations may select different points in and yield distinct attribution patterns even when predictive performance is similar.
@@
EvoXplain does not claim these basins correspond to true causal mechanisms. Rather, it demonstrates that non-identifiability in the training pipeline—including optimisation dynamics, model capacity, and admissible configuration choices—can induce multiple, stable explanatory behaviours.
@@ 5 Scope and Limitations
EvoXplain does not assert that the basins it identifies correspond to true causal mechanisms, nor that particular centroids represent any form of ground truth. Its outputs are contingent upon the attribution method used, and the observed structure may also reflect noise or inductive biases inherent to the model or the explanation technique. EvoXplain does not attempt to resolve these questions. Rather, its purpose is to bring such issues to the forefront and to encourage more cautious interpretation of explanation outputs.
EvoXplain has been designed to assess how models behave across repeated training runs. It determines whether a model settles into a single basin of explanations or whether different explanations emerge across retraining. This perspective comes with several limitations.
@@
SHAP is used here as a widely adopted probe of explanatory structure rather than as a privileged or canonical explanation method; EvoXplain evaluates the stability of attribution geometry under retraining, not the correctness of any particular attribution paradigm.